OpenStudy (anonymous):

identify the conic 5x^2+9y^2-40x-18y+44=0 and give the coordinates of its center

OpenStudy (jdoe0001):

so, the \(x^2 \) and \(y^2\) components are POSITIVE, that's a dead giveaway is an ellipse

OpenStudy (jdoe0001):

well, could be a circle, or an same-axis values ellipse ehehe

OpenStudy (jdoe0001):

the rest is just "completing the square"

OpenStudy (jdoe0001):

do you know how to "complete the square" to get a "perfect square trinomial"?

OpenStudy (anonymous):

I know how to complete the square normally...do you do something special to get a perfect square trinomial?

OpenStudy (anonymous):

Yep, do that 3 times

OpenStudy (jdoe0001):

well, you grab all the "x" elements and the "y" elements and group them together first get a common factor so you're only left inside the parentheses with \(x^2 \ and\ y^2\) as opposed to \(3x^2 \ and \ 5y^2\)

OpenStudy (jdoe0001):

from there, we work with our good jolly friend, Mr 0 [zero]

OpenStudy (anonymous):

ugh I hate completing the square! ok give me a minute

OpenStudy (jdoe0001):

I'll rewrite it some in a few secs

OpenStudy (jdoe0001):

$$ 5x^2+9y^2-40x-18y+44=0\\ (5x^2-40x)+(9y^2-18y)+44\\ \text{so get common factors}\\ 5(x^2-8x+\boxed{?})+9(y^2-2y+\boxed{?})+44=0\\ $$

OpenStudy (jdoe0001):

so, we need to find those " ? " values, and those numbers come from Mr 0 so, if you find 25 and 13, you'd need to SUBSTRACT 25 and 13 because what we're really doing is BORROWING from 0 0 = 25+13-25-13 :)

OpenStudy (anonymous):

wait wait wait. what are 25 and 13?

OpenStudy (anonymous):

where did they come from?

OpenStudy (jdoe0001):

just an example

OpenStudy (anonymous):

oh ok. I was totally lost

OpenStudy (jdoe0001):

just a couple of arbitrary numbers ehhe

OpenStudy (anonymous):

so i need to plug in 0 where the ?'s are?

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

$$ 5x^2+9y^2-40x-18y+44=0\\ (5x^2-40x)+(9y^2-18y)+44\\ \text{so get common factors}\\ 5(x^2-8x+\color{blue}{4^2})+9(y^2-2y+\color{blue}{1^2})+44=0\\ \text{so plus } 5(4^2)+9(1^2), \ \ then \ \ \color{red}{-5(4^2)-9(1^2)}\\ 5(x^2-8x+\color{blue}{4^2})+9(y^2-2y+\color{blue}{1^2}) \color{red}{-5(4^2)-9(1^2)}+44 = 0 $$

OpenStudy (jdoe0001):

so, we're really borrowing from 0, +89, and giving the loan back with -89

OpenStudy (anonymous):

ok what next?

OpenStudy (jdoe0001):

the parenthesized "perfect squares" are just a 2 binomials, so hold on

OpenStudy (jdoe0001):

$$ 5(x-4)^2+9(y-1)^2 \color{red}{-89}+44 = 0\\ 5(x-4)^2+9(y-1)^2 = 45 $$

OpenStudy (jdoe0001):

from there you just make the = 45, to =1 or set the equation to = 1 meaning divide both sides by 45 that is each equation over 45 = 45/45

OpenStudy (anonymous):

ok and that's it?

OpenStudy (jdoe0001):

$$ \cfrac{5(x-4)^2}{45}+ \cfrac{9(y-1)^2}{45}=1\\ \cfrac{(x-4)^2}{9}+ \cfrac{(y-1)^2}{5}=1\\ $$

OpenStudy (jdoe0001):

and that's your ellipse

OpenStudy (anonymous):

ok thank you! you are so helpful

OpenStudy (jdoe0001):

and the coordinates for the center is at (h,k) http://www.mathwarehouse.com/ellipse/images/translations/general_formula_major.gif

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

so center is (-4,-1)

OpenStudy (jdoe0001):

well, (x- (+4)) => (x-4) and (y - (+1) => (y-1)

OpenStudy (jdoe0001):

so, the " - " isn't part of the h,k elements

OpenStudy (anonymous):

so just (4,1)

OpenStudy (jdoe0001):

is part of the formula now if you had (x+4) => (x-(-4)) so the "h" woulld be -4 then

OpenStudy (jdoe0001):

yes, (4,1)